# I see what you mean Co-Speech Gestures for Reference Resolution in Multimodal Dialogue

### v1

@@ Abstract
Figure 1: Example from the CABB dataset Rasenberg et al. (2022), illustrating how participants resolve references through speech and gestures in face-to-face dialogue. The speaker on the right says “there is a circle on the front” while performing a representational gesture. The object is shown for illustration but not visible to the listener; the orange highlight marks the referent as annotated by experts. Our work draws on these interactions to model multimodal reference resolution.
@@ 1 Introduction
However, how to computationally represent and interpret co-speech gestures remains an understudied problem, particularly within Natural Language Processing. Other gestural forms, such as deictic gestures, i.e., pointing Gatt and Paggio (2013); Kennington and Schlangen (2017a); Chen et al. (2021); Kontogiorgos et al. (2018) or beat gestures, i.e., rhythmic movements without semantic content (e.g., Sinclair and Schneider, 2021; Abzaliev et al., 2022), have received some attention. However, the challenges posed by iconic, representational gestures and their contribution to reference resolution have hardly been tackled from a data-driven perspective.
In this paper, we propose an approach to learning embeddings for representational gestures that exploits not only the body movements that make up a gesture, but also the semantically related speech that is typically produced simultaneously with it. We combine contemporary self-supervised learning techniques to train a Transformer-based gesture encoder and ground it in information from features extracted by text or speech large language models. Then, we test the effectiveness of the pre-trained gesture embeddings in the downstream task of reference resolution in face-to-face dialogue, showing that gestures—as learned by our proposed multimodal approach—have significant predictive power that complements the verbal modality. More concretely, we make the following contributions:
@@
  • Our experiments make use of the CABB dataset Rasenberg et al. (2022); Eijk et al. (2022), collected by cognitive scientists. We make available the pre-processed data and the code to reproduce all our results via a public GitHub repository, providing valuable resources to community. [^1] https://github.com/EsamGhaleb/ReferenceResolution
@@ 2 Related Work
Learning multimodal representations.
@@
Reference resolution in dialogue.
Reference resolution in dialogue has mostly been modelled as the task of identifying the referent of text-based linguistic expressions, ignoring non-verbal cues. For example, Skantze and Willemsen (2022) proposed COLLIE, a continual learning method that adjusts language embeddings to accommodate new language use for new referents; in an earlier study Shore and Skantze (2018), the authors found that leveraging dialogue history in the form of previous referring expressions improves model prediction, similarly to Takmaz et al. (2020). Resolving referring linguistic expressions in the visual modality has also been studied in computer vision thanks to datasets such as ReferIt Kazemzadeh et al. (2014), Flicker30k Entities Plummer et al. (2015), and Visual Genome Krishna et al. (2017), which map referring expressions to regions in an image.
@@ 3 Data
We use CABB-S for evaluating our gesture representation pre-training approach (Section 4.3) and for the experiments on reference resolution (Section 5).
@@
Eijk et al. (2022) contains an additional 49 dialogues by 98 different subjects, with about 42,5 hours of recordings. It is therefore much larger than CABB-S. Only 42 dialogues are manually transcribed and no manual annotations regarding gestures are present. To identify gestures, we use the segmentation model by Ghaleb et al. (2024a), resulting in 30k automatically segmented gestures. To increase the amount of data available for pre-training, we oversample by selecting 1-second time windows overlapping more than 50% with the automatically segmented gestures. This results in approximately 400k data samples, which we refer to as CABB-XL. We use Whisper-X Bain et al. (2023) to automatically generate speech transcriptions when manual transcriptions are unavailable. 83% of the gestures are accompanied by speech. We use CABB-L and CABB-XL for pre-training the models introduced in Section 4.
@@
To process body movements, we apply the procedure used by Ghaleb et al. (2024b) to CABB-S, CABB-L, and CABB-XL. Concretely, we sample 1-second time windows centered around each segmented gesture and use MMPose Sengupta et al. (2020) to extract skeletal information, i.e., 2D keypoint coordinates for 27 upper body and hand joints. To process the verbal modality, we extract 2-second windows centered around the sampled gestures to account for the temporal asynchrony between speech and gestures Holler and Levinson (2019a), and use both the raw speech and the transcriptions as described in the next section.
@@ 4.1 Modality Encoders
Although wav2vec-2 representations may capture diverse linguistic properties including prosody, phonetics, and to some extent semantics Tsai et al. (2022); Zaiem et al. (2025), they are less semantically rich than word embeddings learned from text. Therefore, we also experiment with the word embeddings from a pre-trained Dutch BERT-based model de Vries et al. (2019).
@@ 4.2 Model Architectures
We propose three pre-training strategies to learn gesture representations. The first one is unimodal, in the sense that it learns representations only considering the body movements that make up a gesture. The other two (multimodal and multimodal-x) are motivated by a more holistic conception of co-speech gestures as multimodal acts Holler and Levinson (2019b); Özyürek (2014), and therefore exploit both skeletal and concurrent verbal input. We describe the gist of each architecture here and provide further technical details in Appendix B.
@@
This model jointly optimizes a masked reconstruction loss and a unimodal contrastive loss. For the former, we follow the original procedure for pre-training DSTFormer Zhu et al. (2023) by randomly masking portions of the 2D keypoint skeletal input and learning to reconstruct them. The unimodal contrastive loss pulls representations of two views of augmented skeletal data closer while pushing them away from other negative samples in a batch. A detailed diagram of this architecture is shown in Figure 10, Appendix B.
@@
This model combines the two losses from the unimodal architecture with a multimodal contrastive loss that integrates information from skeletons and either speech or semantics. For the latter, we use a CLIP-like contrastive objective Radford et al. (2021) mapping global representations of gestures and co-occurring utterances (as either raw speech or semantics) in a joint feature space.
@@
This approach is optimized by a combination of two contrastive losses as shown in Figure 2. The first (blue line in the figure) aligns skeletal embeddings with global semantic representations. The second loss (green line), in turn, aligns unimodal gesture representations with fused gesture–crossmodal embeddings. These embeddings are obtained by ingesting semantic information (e.g. text tokens) into our adapted DSTFormer via cross-attention blocks, previously introduced in Section 4.1.
@@
We train the three architectures above using CABB-L and CABB-XL, which allows us to test the impact of increasing the size of the training data. In Appendix C, we provide further details about the backbone models, projection heads, and the parameters used in the learning objectives, along with the implementation details.
@@ 4.3 Evaluation
We compare our models against the framework by Ghaleb et al. (2024b). This work proposed a pre-training approach to gesture representation learning based on Spatio-Temporal Graph Convolutional Networks (ST-GCN) using unimodal and multimodal contrastive learning with co-occurring raw speech. For comparability with our approach, when reproducing this framework, in addition to raw speech we extend it to also use text-based semantics and pre-train it on CABB-XL. [^3] Due to the architecture of ST-GCN, it is not possible to combine it with the multimodal-x architecture introduced in our work (Figure 2). The comparison of correlation coefficients for form similarity is shown in Figure 4. As can be observed, the gesture representations learned by our transformer-based encoder are more aligned with form-based expert annotations as evidenced by higher correlation values across the board.
@@ 5 Reference Resolution
More concretely, given a gesture unseen during pre-training, we investigate two scenarios: (1) We measure how accurate a reference resolution system that only has access to the gesture (i.e., to skeletal input) is at predicting the gesture’s referent. Here the gesture embedding is extracted zero-shot with our models, some of which exploited raw speech or text semantics during pre-training—but importantly verbal input is not provided at inference time in this scenario. (2) In the second scenario, at inference time the reference resolution system has access to both the unseen gesture to be resolved and any concurrent speech. We operationalise this by concatenating the gesture embedding extracted with our models and a semantic embedding of the concurrent speech, and then measure whether this leads to higher reference resolution accuracy than only exploiting the semantics of concurrent speech.
Resolution model and evaluation setup
The resolution model leverages our model architectures pre-trained on CABB-XL without any fine-tuning. The model is implemented as a multi-class MLP classifier with two hidden layers of size 300 and 150, respectively, and it is trained on CABB-S. Given a gesture encoded with our pre-trained models, we train the MLP to predict one referent among 70 possible object sub-parts (see Appendix A for details) using a batch size of 32 and a learning rate of $10^{-4}$ with the Adam optimizer for 200 epochs. Recall from Section 3 that each dialogue consists of six rounds. We use leave-one-round-out cross-validation, holding out the gestures in one round as a test set and training on the gestures in the remaining rounds across all dialogues in CABB-S. We use accuracy as an evaluation metric. Chance accuracy is approximately 1.4%, and a model using random gesture embeddings (without access to our pre-trained models) achieves around 3% accuracy.
@@ 5.1 Gesture-Only Reference Resolution
Figure 5: Average reference resolution accuracies for gesture-only embeddings are shown for unimodal, multimodal, and multimodal-x representations. The multimodal and multimodal-x models are pre-trained with text-based semantic input, and the embeddings are derived only from skeletal data.
We first evaluate the resolution model when it only has access to gestural information (i.e., skeletal data) as input. As shown in Figure 5, when a gesture is encoded with our unimodal model, the average resolution accuracy is 16%, significantly above the random baselines. Using embeddings from models that were pre-trained jointly with text-based semantics significantly increases resolution accuracy to around 19%, with no statistically significant difference between the multimodal and multimodal-x approaches. These results show that our pre-trained gesture representations capture information that is useful to identify referents. Moreover, they indicate that learning gesture representations by jointly exploiting body movements and the semantics of co-occurring speech enhances their reference resolution potential, even when information about concurrent speech is not provided at prediction time.
Figure 6: Average accuracies for reference resolution using gestures and co-occurring verbal information are reported for unimodal gesture embeddings, multimodal-x gesture embeddings, semantic embeddings, and their concatenated representations. The multimodal-x gesture embeddings are learned through pre-training with semantic embeddings.
@@ 5.2 Reference Resolution with Gestural and Co-occurring Verbal Information
Next, we assess the resolution model when it has access to gestures and the speech co-occurring with them. speech We operationalise this scenario by concatenating gesture embedding (extracted with either unimodal or multimodal-x pre-trained models) [^4] Given the lack of statistically significant difference between multimodal and multimodal-x in Section 5.1 Fig. 5, we focus on the multimodal-x model for this experiment. with a semantic embedding derived from the transcribed co-occurring speech using BERTje de Vries et al. (2019). The results are shown in Figure 6, where we also include a condition where the reference resolution model exclusively uses the concurrent speech in the form of a semantic embedding. In that condition, resolution accuracy is 24%. That is, the concurrent information present in the verbal modality has stronger predictive power to identify referents than body movements alone, which is to be expected in spoken conversations. Importantly, when both the verbal and gestural modalities are combined, we observe a significant increase in reference resolution accuracy, reaching 31% when gestures are encoded with our multimodal-x model, pre-trained with text-based semantics. These findings confirm the complementary roles of gesture and speech in reference resolution and highlight the benefits of exploiting such complementarity for gesture representation learning.
@@ 5.3 Impact of Dialogue History
To isolate the impact of dialogue history on gesture-driven reference resolution, in this experiment we focus on identifying referents with only gestural information as input (as in Section 5.1), [^5] The impact of dialogue history on text-based reference resolution has already been extensively studied, e.g., Shore and Skantze (2018); Haber et al. (2019); Takmaz et al. (2020); Hawkins et al. (2020). comparing our unimodal and multimodal-x pre-trained models. Figure 7 shows that as the conversation unfolds over the rounds, the dialogue-specific reference resolution models clearly outperform the baselines (dotted lines). Our statistical analysis shows that there is a significant difference in accuracy between the two (independent t-test yielding $t=2.9$, $p\ll 0.05$ for both the unimodal and multimodal-x models). It also indicates that the pattern of increased accuracy, as more dialogue history becomes available, is more pronounced when the gestures are encoded with the multimodal-x pre-trained model (Spearman correlation between accuracy values and dialogue round numbers: $\rho=0.32$ for the unimodal model and $\rho=0.35$ for the multimodal-x model, with $p\ll 0.05$ in both instances). [^6] Note that there is no statistically significant difference in accuracy between rounds 5 and 6, despite the apparent drop.

### v3

@@ Abstract
Figure 1: Example from the CABB dataset Rasenberg et al. (2022), illustrating how participants resolve references through speech and gestures in face-to-face dialogue. The speaker on the right says “there is a circle on the front” while performing a representational gesture that resembles a circle. The object discussed is shown for illustration but not visible to the listener; the orange arrow points to the referent as annotated by experts. Our work draws on these interactions to model multimodal reference resolution.
@@ 1 Introduction
However, how to computationally represent and interpret co-speech gestures remains an understudied problem, particularly within Natural Language Processing. Other gestural forms, such as deictic gestures, i.e., pointing Gatt and Paggio (2013); Kennington and Schlangen (2017a); Chen et al. (2021); Kontogiorgos et al. (2018) or beat gestures, i.e., rhythmic movements without semantic content (e.g., Sinclair and Schneider, 2021; Abzaliev et al., 2022), have received some attention. In contrast, the challenges posed by iconic, representational gestures and their contribution to reference resolution have hardly been tackled from a data-driven perspective.
In this paper, we propose an approach to learning embeddings for representational gestures that exploits not only the body movements that make up a gesture but also the semantically related speech that is typically produced simultaneously with it. We combine contemporary self-supervised learning techniques to train a Transformer-based gesture encoder and ground it in information from features extracted by text or speech large language models. Then, we test the effectiveness of the pre-trained gesture embeddings in the downstream task of reference resolution in face-to-face dialogue, showing that gestures—as learned by our proposed multimodal approach—have significant predictive power that complements the verbal modality. More concretely, we make the following contributions:
@@
  • Our experiments make use of the CABB dataset Rasenberg et al. (2022); Eijk et al. (2022), collected by cognitive scientists. We make available the pre-processed data and the code to reproduce all our results via a public GitHub repository, providing valuable resources to the community. [^1] https://github.com/EsamGhaleb/ReferenceResolution
@@ 2 Related Work
2.1 Learning Multimodal Representations
@@
2.2 Reference Resolution in Dialogue
Reference resolution in dialogue has mostly been modelled as the task of identifying the referent of text-based linguistic expressions, ignoring non-verbal cues. For example, Skantze and Willemsen (2022) proposed COLLIE, a continual learning method that adjusts language embeddings to accommodate new language use for new referents; in an earlier study, Shore and Skantze (2018) found that leveraging dialogue history in the form of previous referring expressions improves model prediction, similarly to Takmaz et al. (2020). Resolving linguistic referring expressions in the visual modality has also been studied in the field of computer vision thanks to datasets such as ReferIt Kazemzadeh et al. (2014), Flicker30k Entities Plummer et al. (2015), and Visual Genome Krishna et al. (2017), which map referring expressions to regions in an image.
@@ 3 Data
We use CABB-S for evaluating our pre-training approach to gesture representation learning (Section 4.3) and for the experiments on reference resolution (Section 5).
@@
Eijk et al. (2022) contains an additional 49 dialogues by 98 different subjects, with about 42.5 hours of recordings. It is therefore much larger than CABB-S. Only 42 dialogues are manually transcribed and no manual annotations regarding gestures are present. To identify gestures, we use the segmentation model by Ghaleb et al. (2024a), which has been shown to achieve a mean Average Precision (mAP) of 76% on the CABB-S dataset. Applying this model to CABB-L results in 30k automatically segmented gestures. [^3] Some qualitative results on segmentation performance can be found in Appendix B.
To increase the amount of data available for pre-training, we oversample by selecting 1-second time windows overlapping more than 50% with the automatically segmented gestures. This results in approximately 400k data samples, which we refer to as CABB-XL. We use Whisper-X Bain et al. (2023) to automatically generate speech transcriptions when manual transcriptions are unavailable. 83% of the gestures are accompanied by speech. We use CABB-L and CABB-XL for pre-training the models introduced in Section 4.
@@
To process body movements, we apply the procedure used by Ghaleb et al. (2024b) to CABB-S, CABB-L, and CABB-XL. Concretely, we sample 1-second time windows centered around each segmented gesture and use ViTPose Xu et al. (2022) to extract skeletal information, i.e., 2D keypoint coordinates for 27 upper body and hand joints. Holler and Levinson (2019a) showed that related speech is often produced before or after gesture strokes (the most meaning-bearing segment of a gesture) by a few hundred milliseconds. To account for this temporal asynchrony between speech and gestures, when processing the verbal modality we extract 2-second windows centered around the sampled gestures, and use both the raw speech and the transcriptions as described in the next section.
@@ 4.1 Modality Encoders
Although wav2vec-2 representations may capture diverse linguistic properties including prosody, phonetics, and to some extent semantics Tsai et al. (2022); Zaiem et al. (2025), they are less semantically rich than word embeddings learned from text. Therefore, we also experiment with the word embeddings from a pre-trained Dutch BERT-based model (BERTje; de Vries et al., 2019).
@@ 4.2 Model Architectures
We propose three pre-training strategies to learn gesture representations. The first one is unimodal, in the sense that it learns representations only considering the body movements that make up a gesture. The other two (multimodal and multimodal-X) are motivated by a more holistic conception of co-speech gestures as multimodal acts Holler and Levinson (2019b); Özyürek (2014), and therefore exploit both skeletal and concurrent verbal input. We describe the gist of each architecture here and provide further technical details in Appendix C.
@@
This model only leverages skeletal information. It jointly optimizes a masked reconstruction loss and a unimodal contrastive loss. For the former, we follow the original procedure for pre-training DSTFormer Zhu et al. (2023) by randomly masking portions of the 2D keypoint skeletal input and learning to reconstruct them. The unimodal contrastive loss pulls representations of two views of augmented skeletal data closer while pushing them away from other negative samples in a batch. A detailed diagram of this architecture is shown in Figure 14, Appendix C.
@@
This model combines the two losses from the unimodal architecture with a multimodal contrastive loss that integrates skeletal information with either speech or semantics. For the multimodal contrastive loss, we use a CLIP-like contrastive objective Radford et al. (2021) mapping global representations of gestures and co-occurring utterances (as either raw speech or semantics) into a joint feature space.
@@
Finally, this model is optimized with two complementary contrastive losses, as illustrated in Figure 2: the multimodal contrastive loss (blue arrows) described as part of the multimodal architecture above and an additional crossmodal contrastive loss (green arrows), which aligns the unimodal skeleton representation with the fused gesture-semantic (or gesture-speech) embedding—the latter is obtained by injecting text tokens (or speech frames) into our DSTFormer backbone via the cross-attention layers described in Section 4.1.
The unimodal and multimodal architectures can thus be seen as ablations of the arguably more powerful multimodal-X architecture: the multimodal architecture does not include cross-attention layers, and the unimodal architecture omits multimodal alignment altogether.
@@
We train the three architectures above using CABB-L and CABB-XL, which allows us to test the impact of increasing the size of the training data. In Appendix D, we provide further details about the backbone models, projection heads, and the parameters used in the learning objectives, along with the implementation details.
@@ 4.3 Evaluation
We compare our models against the framework by Ghaleb et al. (2024b). This work proposed a pre-training approach to gesture representation learning based on Spatio-Temporal Graph Convolutional Networks (ST-GCN) using unimodal and multimodal contrastive learning with co-occurring raw speech. For comparability with our approach, when reproducing this framework, in addition to raw speech we extend it to also use text-based semantics and pre-train it on CABB-XL. [^4] Due to the architecture of ST-GCN, it is not possible to combine it with the multimodal-X architecture introduced in our work (Figure 2). The comparison of correlation coefficients with form similarity is shown in Figure 4. As can be observed, the gesture representations learned by our Transformer-based encoder are more aligned with form-based expert annotations as evidenced by higher correlation values across the board.
@@
Figure 5: Gesture-based reference resolution. The reference resolution classifier leverages gesture representations encoded with our pre-trained models.
@@ 5 Reference Resolution
5.1 Resolution Model and Evaluation Setup
The resolution model leverages our model architectures pre-trained on CABB-XL without any fine-tuning. This is schematically shown in Figure 5. The model is implemented as a multi-class MLP classifier with two hidden layers of size 300 and 150, respectively, and it is trained on CABB-S. Given a gesture encoded with our pre-trained models, we train the MLP to predict one referent among 70 possible object sub-parts (see Appendix A for details) using a batch size of 32 and a learning rate of $10^{-4}$ with the Adam optimizer for 200 epochs. Recall from Section 3 that each dialogue consists of six rounds. We use leave-one-round-out cross-validation, holding out the gestures in one round as a test set and training on the gestures in the remaining rounds across all dialogues in CABB-S. We use accuracy as an evaluation metric. Since speakers can refer to any of 70 possible sub-parts (across the 16 objects), randomly guessing results in a rate of 1/70, leading to approximately 1.4% accuracy. A model using random gesture embeddings (without access to our pre-trained models) achieves around 3% accuracy.
Given a gesture unseen during pre-training, we investigate two scenarios: In Section 5.2, we measure how accurate a reference resolution system that only has access to the gesture (i.e., to skeletal input) is at predicting the gesture’s referent. Here the gesture embedding is extracted zero-shot with our models, some of which exploited raw speech or text semantics during pre-training—but importantly verbal input is not provided at inference time in this scenario. In Section 5.3, we consider a second scenario, where at inference time the reference resolution system has access to both the unseen gesture to be resolved and any concurrent speech. We operationalise this by concatenating the gesture embedding extracted with our models and a semantic embedding of the concurrent speech, and then measure whether this leads to higher reference resolution accuracy than only exploiting the semantics of concurrent speech.
@@ 5.2 Gesture-Only Reference Resolution
Figure 6: Average reference resolution accuracies for gesture-only embeddings are shown for unimodal, multimodal, and multimodal-x representations. The multimodal and multimodal-x models are pre-trained with text-based semantic input, and the embeddings are derived only from skeletal data. Asterisks indicate statistically significant differences, with ** $p<0.01$ and *** $p<0.001$; ‘ns’ stands for ‘not significant’.
We first evaluate the resolution model when it only has access to gestural information (i.e., skeletal data) as input. As shown in Figure 6, when a gesture is encoded with our unimodal model, the average resolution accuracy is 16%, significantly above the random baselines. Using embeddings from models that were pre-trained jointly with text-based semantics significantly increases resolution accuracy to around 19%, with no statistically significant difference between the multimodal and multimodal-X approaches. These results show that our pre-trained gesture representations capture information that is useful to identify referents. [^5] In Appendix  E, we present a supplementary experiment showing that this is the case even in the presence of noisy skeletal data, which further emphasizes the robustness of our gesture representations. Moreover, they indicate that learning gesture representations by jointly exploiting body movements and the semantics of co-occurring speech enhances their reference resolution potential, even when information about concurrent speech is not provided at prediction time.
Figure 7: Average accuracies for reference resolution using gestures and co-occurring verbal information are reported for unimodal gesture embeddings, multimodal-x gesture embeddings, semantic embeddings, and their concatenated representations. The multimodal-x gesture embeddings are learned through pre-training with semantic embeddings. Asterisks indicate statistically significant differences, with * $p<0.05$, ** $p<0.01$ and *** $p<0.001$.
@@ 5.3 Reference Resolution with Gestural and Co-occurring Verbal Information
Next, we assess the resolution model when it has access to gestures and the speech co-occurring with them. As mentioned above, we operationalise this scenario by concatenating a gesture embedding (extracted with either unimodal or multimodal-X pre-trained models) [^6] Given the lack of statistically significant difference between multimodal and multimodal-X in Section 5.2 Fig. 6, we focus on the multimodal-X model for this experiment. with a semantic embedding derived from the transcribed co-occurring speech using BERTje de Vries et al. (2019). The results are shown in Figure 7, where we also include a condition where the reference resolution model exclusively uses the concurrent speech in the form of a semantic embedding. In that condition, resolution accuracy is 24%. That is, the concurrent information present in the verbal modality has stronger predictive power to identify referents than body movements alone, which is to be expected in spoken conversations. Importantly, when both the verbal and gestural modalities are combined, we observe a significant increase in reference resolution accuracy, reaching 31% when gestures are encoded with our multimodal-X model, pre-trained with text-based semantics. These findings confirm the complementary roles of gesture and speech in reference resolution and highlight the benefits of exploiting such complementarity for gesture representation learning.
@@ 5.4 Impact of Dialogue History
To isolate the impact of dialogue history on gesture-driven reference resolution, in this experiment we focus on identifying referents with only gestural information as input (as in Section 5.2), [^7] The impact of dialogue history on text-based reference resolution has already been extensively studied, e.g., Shore and Skantze (2018); Haber et al. (2019); Takmaz et al. (2020); Hawkins et al. (2020). comparing our unimodal and multimodal-x pre-trained models. Figure 8 shows that as the conversation unfolds over the rounds, the dialogue-specific reference resolution models outperform the baselines (dotted lines). Our statistical analysis shows that there is a significant difference in accuracy between the two (independent t-test yielding $t=2.9$, $p\ll 0.05$ for both the unimodal and multimodal-x models). It also indicates that the pattern of increased accuracy, as more dialogue history becomes available, is more pronounced when the gestures are encoded with the multimodal-x pre-trained model (Spearman correlation between accuracy values and dialogue round numbers: $\rho=0.32$ for the unimodal model and $\rho=0.35$ for the multimodal-x model, with $p\ll 0.05$ in both instances). [^8] Note that there is no statistically significant difference in accuracy between rounds 5 and 6, despite the apparent drop.